Nomzi Traders sold furniture to a customer for R1 610 cash, the mark up on cost is 20%. Assuming Nomzi Traders uses a perpetual inventory system, is a registered VAT vendor and the VAT rate is 15%. (Round your answer off to the nearest Rand.) 

What will the cost of sales be for this transaction?

 


1
Expert's answer
2021-06-24T12:05:33-0400

Solution:

Sales = Cost + mark-up

When calculating cost and sale price, VAT is normally excluded.

Therefore, we start by first eliminating VAT from the sales:

Sales inclusive of 15"\%" VAT = 1,610


Sales exclusive of 15"\%" VAT = "\frac{100}{115} \times 1,610 = 1,400"


Now, determine cost:

Cost + mark-up = Sales

100"\%" + 20"\%" = 120"\%"

100"\%" + 20"\%" = 1,400


Therefore:

Cost of sales = "\frac{100}{120} \times 1,400 = 1,166.67 = 1,167"


The cost of sales for this transaction = R1,167